# Mimicking Human Intuition: Cognitive Belief-Driven Reinforcement Learning

### v3

@@ Abstract
Traditional reinforcement learning (RL) methods mainly rely on trial-and-error exploration, often lacking mechanisms to guide agents toward more informative decision-making and struggling to leverage past experiences, resulting in low sample efficiency. To overcome this issue, we propose an innovative framework inspired by cognitive principles: Cognitive Belief-Driven Reinforcement Learning (CBD-RL). By incorporating cognitive heuristics, CBD-RL transforms conventional trial-and-error learning into a more structured and guided learning paradigm, simulating the human reasoning process. This framework’s core is a belief system that optimizes action probabilities by integrating feedback with prior experience, thus enhancing decision making under uncertainty. It also organizes state-action pairs into meaningful categories, promoting generalization and improving sample efficiency. The concrete implementations of this framework, CBDQ, CBDPPO, and CBDSAC, demonstrate superior performance in discrete and continuous action spaces in diverse environments such as Atari and MuJoCo. By bridging cognitive science and reinforcement learning, this research opens a new avenue for developing RL systems that are more interpretable, efficient, and cognitively inspired.
Machine Learning, ICML
@@ 1 Introduction
Figure 1: Inspired by the way pets intuitively choose to walk, stand, or jump in different settings (path, ocean, river)
Reinforcement learning (RL) has achieved strong performance in domains with well-defined structure, such as Atari (Mnih et al., 2015) and continuous control (Lillicrap, 2015; Qiao et al., 2024a), yet remains markedly less sample-efficient than human learning. Humans can generalize from limited experience by abstracting structural regularities and transferring them across contexts, while RL agents typically require millions of interactions to attain basic competence. This gap reflects the absence of cognitive inductive biases crucial to efficient generalization—namely, the integration of uncertain evidence with prior expectations (Griffiths & Tenenbaum, 2005; Peterson & Beach, 1967), and the formation of reusable abstractions that support compositional reasoning (Tenenbaum et al., 2011; Lake et al., 2015).
Cognitive science highlights belief updating and conceptual abstraction as core mechanisms enabling efficient human learning. Probabilistic belief revision, formalized via Bayesian inference, allows humans to construct generative models for inference under uncertainty (Gigerenzer et al., 1991; Tenenbaum et al., 2006). In parallel, conceptual abstraction compresses experience into structured forms—such as taxonomies and causal schemas—that support generalization and reuse (Tenenbaum & Griffiths, 2001; Gopnik & Wellman, 2012). Together, these mechanisms induce compact inductive biases that facilitate data-efficient decision-making, in contrast to model-free RL, which lacks structured representations of environment dynamics (Lake et al., 2017; Botvinick & Weinstein, 2014).
While recent RL research has begun to emulate facets of human flexibility, its progress remains fragmented. Bayesian and model-based approaches propagate uncertainty (Ghavamzadeh et al., 2015), but fail to integrate concepts across contexts. Multi-policy fusion reuses pretrained skills (Chiu et al., 2023; Peng et al., 2021), although static libraries hinder online abstraction. Each line of work addresses a single aspect—temporal hierarchy, uncertainty reasoning, or skill reuse—yet none unifies probabilistic inference with dynamic concept formation throughout learning. This fragmentation reflects a deeper epistemological shift. Sutton and Silver advocate for an “Era of Experience,” where intelligence arises not from data-driven prediction, but from the agent’s active organization of its own interaction history(Silver & Sutton, 2025). Experience, in this view, is the substrate of knowledge—concepts and strategies emerge through continual abstraction from behavior.
Based on this proactive insights, we propose Cognitive Belief‑Driven Reinforcement Learning (CBD‑RL), a unified framework that integrates probabilistic inference and conceptual abstraction into the RL process. At its core, a Smoothed Bellman Operator replaces pointwise maximization with belief-weighted expectations, enabling uncertainty-aware value updates. A Conceptual Category Formation(CCF) module clusters state–action pairs into latent abstractions, supporting modular reuse and compositional generalization. We implement CBD-RL in Q-learning(Watkins et al., 1992), SAC (Haarnoja et al., 2018), and PPO (Schulman et al., 2017), and show consistent improvements in sample efficiency, final return, and generalization across both discrete and continuous domains. We hope we can connect cognitive science in an innovative and unique way to enhance the effectiveness of RL algorithms.
@@ 2 Related Works
2.1 Value-based and Policy-based Reinforcement Learning Methods
Model-free reinforcement learning (RL), including value-based methods like XQL (Garg et al., 2023) and DoubleGum (Hui et al., 2023), and policy-based approaches such as PPO (Schulman et al., 2017), has seen improved stability and exploration through stochastic regularization and adaptive experience replay (Hassani et al., 2025). However, these methods often fail to retain and reuse knowledge effectively, especially in continual learning settings (Dohare et al., 2024). Replay strategies—hierarchical (Yin & Pan, 2017), sequence-based (Li et al., 2024), and selective (De Bruin et al., 2018)—focus on raw data reuse without abstraction, limiting conceptual generalization and structure-aware transfer (Jeen et al., 2023).
2.2 Cognitive Science Perspectives on Efficient Learning and Decision-Making
@@
Humans exhibit exceptional cognitive efficiency, generalizing from limited experience through Bayesian inference that integrates prior knowledge with new evidence under uncertainty (Tenenbaum & Griffiths, 2001; Griffiths & Tenenbaum, 2005; Tenenbaum et al., 2006). This process supports conceptual abstraction—extracting high-level structure from sparse data—and enables causal reasoning and cross-domain transfer (Tenenbaum et al., 2011; Kemp & Tenenbaum, 2008). Recent studies formalize how human learners reorganize internal knowledge via probabilistic reasoning (Lake et al., 2015, 2017), motivating the integration of such cognitive principles into machine learning to enhance scalability, adaptability, and sample efficiency (Ma et al., 2022). Other work shows that uncovering latent causal structures—even in domains like joint behavior—can enhance model interpretability and abstraction (Gu et al., 2024a, b).
2.3 Trade‑Offs in Abstraction and Policy Integration Paradigms
Reinforcement-learning agents adopt diverse mechanisms for abstraction and policy reuse, each with characteristic trade-offs. Hierarchical methods, such as Option-Critic (Bacon et al., 2017) and FeUdal Networks (Vezhnevets et al., 2017), acquire temporally extended skills but struggle to adjust option granularity and time-scales on the fly. Knowledge-grounded approaches (Jiang & Luo, 2019; Kimura et al., 2021) embed symbolic or relational priors—often via graph or logic modules—improving sample efficiency at the cost of external knowledge engineering. Curriculum-style transfer and progressive networks (Rusu et al., 2016; Narvekar et al., 2020) accelerate learning by reusing feature hierarchies, yet require carefully staged task sequences and exhibit limited structural flexibility. Multi-policy fusion techniques (Qiao et al., 2024b; Peng et al., 2021; Chiu et al., 2023) blend diverse pre-trained controllers through adversarial or attention mechanisms, typically assuming a fixed policy library that hampers online abstraction. These paradigms expose a common tension between structural rigidity, prior dependence, and scalability—motivating alternatives that infer latent concepts and reason about uncertainty directly from experience.
@@ 3 Problem Formulation
Markov Decision Processes (MDP) To solve an RL problem, the agent optimizes the control policy under an MDP $\mathcal{M}$, which can be defined by a tuple $(\mathcal{S},\mathcal{A},T,r,\mu_{0},\gamma,T)$ where: 1) $\mathcal{S}$ and $\mathcal{A}$ denote the space of states and actions. 2) $\mathcal{T}(s_{t+1}|s_{t},a_{t})$ and $r(s_{t},a_{t})$ define the transition probability and reward function. 3) $\mu_{0}$ defines the initial state distribution. 4) $\gamma\in(0,1)$ is the discount factor and $T$ defines the planning horizon. The goal of the RL policy $\pi$ is to maximize expected discounted rewards:
$\arg\max_{\pi}\mathbb{E}_{\pi,\mathcal{T},\mu_{0}}\Big{[}\sum_{t=0}^{T}\gamma^{t}r(s_{t},a_{t})\Big{]}$ (1)
We define the action value function given a policy $\pi$:
$Q(s,a)=\mathbb{E}_{\pi,\mathcal{T},\mu_{0}}\big{[}\sum_{t=0}^{T}\gamma^{t}r(s_{t},a_{t})\mid s_{0}=s,a_{0}=a\big{]}$ (2)
and value function is:
$V(s)=\mathbb{E}_{\pi,\mathcal{T},\mu_{0}}\big{[}\sum_{t=0}^{T}\gamma^{t}r(s_{t},a_{t})\mid s_{0}=s\big{]}$ (3)
4 Methodology
In this section, we introduce the Cognitive Belief-Driven Reinforcement Learning (CBD-RL) framework, which enhances the utilization and decision making of experience by integrating cognitive principles. It incorporates the Smoothed Bellman Operator for probabilistic reasoning (See in Section 4.1) and organizes experiences into conceptual categories through conceptualized experience abstraction, thereby improving learning efficiency and adaptability in complex environments.
4.1 Smoothed Bellman Operator
Human reward processing operates under probabilistic principles, with dopamine neurons encoding both the magnitude and uncertainty of rewards through phasic responses (Dayan & Daw, 2008; Schultz et al., 1997). Rather than responding to fixed values, humans evaluate rewards based on perceived reliability, enabling adaptive decision-making under uncertainty (Schultz, 2015; O’Doherty et al., 2004). These cognitive mechanisms align with behavioral economic models of subjective utility and provide a biological foundation for incorporating probabilistic reasoning into reinforcement learning frameworks (Sutton & Barto, 2018). To align reinforcement learning with human-like probabilistic decision-making, we revisit the classical Bellman operator, which updates value functions deterministically, $\mathcal{T}(s,a)=r_{t}+\gamma\max_{a}Q_{t}(s_{t+1},a).$ The use of $\max$ implies deterministic action selection, neglecting uncertainty and preference variability. In contrast, human decisions are often modeled by Subjective Expected Utility Theory (SEUT) (Mongin & Philippe, 1998; Johnson-Laird et al., 2015), which integrates both reward expectations and belief-driven preferences. Inspired by SEUT, we introduce a probabilistic evaluation mechanism through the distribution $q_{t}(a\mid s_{t+1})$, representing the likelihood of selecting action $a$ given state $s_{t+1}$. This formulation captures the inherent uncertainty in action selection and is formalized in the following proposition.
Proposition 4.1.
Consider a decision-making scenario within an MDP. Let $q_{t}(a\mid s_{t+1})$ represent the probability distribution over actions at the next state $s_{t+1}$, derived from the reward signal. Then, the expected utility $U_{t}(s_{t},x)$ at time $t$ is given by:
$U_{t}(s_{t},x)=\sum_{a\in A}q_{t}(a\mid s_{t+1})u_{t}(s_{t},x),$ (4)
where $u_{t}(s_{t},x)$ denotes the utility of outcome $x$ in state $s_{t}$.
The above proposition formalizes expected utility in reinforcement learning by combining probabilistic evaluations and utility functions, providing a framework for incorporating uncertainty and preference variability into decision-making. To address the limitations of the deterministic $\max$ operation, we propose the Smoothed Bellman Operator, which replaces $\max$ with a probability distribution $q_{t}(a\mid s_{t+1})$, enabling probability-weighted action selection. This smoothing strategy is further detailed in the Appendix B, including formulations such as softmax smoothing to enhance stability and performance in reinforcement learning. Theoretical properties of the Smoothed Bellman Operator, including a Jensen-type inequality and convergence conditions, are provided in Appendix B.
Lemma 4.2 (Convergence of Smoothed Bellman Operator).
$\lim_{t\to\infty}\max_{a}q_{t}(a\mid s_{t+1})=1,$ (5)
4.2 Cognitive Belief-Driven Reinforcement Learning (CBD-RL) Framework
Algorithm 1 Cognitive Belief-Driven RL: Experience Integration via Conceptual Structures
0:
1:  Conceptual Category Set $\{C_{n}\}_{n=1}^{N}$
2:  Adaptive Integration Coefficient $\beta_{t}\in[0,1]$
3:  for Each Category $C_{k}$ do
4:     Initialize Belief Model $P_{k}(a|s)$;
5:  end for
6:  for Each Training Step $t$ do
7:     Observe current state $s_{t}$, determine category $C_{k}$;
8:     Collect reward signal $r_{t}$, environment transition $s_{t+1}$;
9:     Integrate action and prior belief into updated preference:
$b_{t}(a|s_{t})=(1-\beta_{t})\cdot\mathcal{Z}_{t}(a|s_{t+1})+\beta_{t}\cdot P_{k}(a|s_{t})$
10:     Select action $a_{t}\sim b_{t}(\cdot|s_{t})$, interact with environment;
11:     Update $P_{k}$ based on observed outcome and adaptation rule.
12:  end for
In the previous section, we introduce the Smoothed Bellman Operator to incorporate probabilistic evaluation of rewards, drawing inspiration from human cognitive processing. Cognitive science shows that humans efficiently abstract concepts from limited experience and apply them to novel contexts by recognizing structural similarities and shared features (Tenenbaum et al., 2011; Lake et al., 2015; Kemp & Tenenbaum, 2008). Conceptual organization supports memory compression (Gershman et al., 2015), rapid generalization (Tenenbaum et al., 2006), and cross-scenario transfer (Botvinick et al., 2019). Building on these insights, we propose the Cognitive Belief-Driven Reinforcement Learning (CBD-RL) framework, which clusters similar state–action pairs into conceptual categories represented by probability distributions to capture uncertainty and integrate reward signals. This abstraction mechanism enhances experience efficiency and generalization in reinforcement learning.
CBD-RL centers on Conceptual Category Formation (CCF) (Definition 4.4) (Rosch, 1978; Murphy, 2004), which clusters state–action experiences into semantically coherent groups. Each category is parameterized by a Belief distribution (Definition 6) (Premack & Woodruff, 1978; Dennett, 1988), representing probabilistic action preferences within its context. Cognitive science indicates that such abstraction and uncertainty-aware reasoning jointly support generalization from limited data and robust decision-making (Tenenbaum & Griffiths, 2001; Griffiths et al., 2010; Lake et al., 2017; Rogers & McClelland, 2004).
Definition 4.3 (Belief).
A belief represents an agent’s internal representation and understanding of the decision-making context, particularly over the action space. Unlike traditional Bayesian cognitive models, which define belief as a posterior probability distribution, we propose a more direct representation of belief grounded in probability theory (Gigerenzer et al., 1991; Griffiths et al., 2010).
For each state s, we define a belief as a probability measure $\mathcal{P}_{k}$ over the action space $A$, satisfying the following conditions:
[…]
Definition 4.4 (Conceptual Category Formation).
A Conceptual Category Formation (CCF) is a partition of the state space $S$ into meaningful subsets, enabling an agent to generalize from similar experiences. Facilitates efficient learning by grouping states with shared characteristics, promoting rapid adaptation and knowledge transfer. Formally, a CCF of $S$ is a finite partition $C=\{C_{1},\dots,C_{N}\}$ that satisfies the following conditions:
  1. 1.
    Completeness and Exclusivity: The partition covers the entire state space without overlap, i.e., $S=\bigcup_{n=1}^{N}C_{n}$, with $C_{i}\cap C_{j}=\emptyset$ for $i\neq j$.
  2. 2.
    Semantic Coherence: Each category $C_{n}$ is characterized by a representative state $c^{i}_{n}\in C_{n}$, such that for all $c^{j}_{n}\in C_{n}$ $i\neq j$, $d(c^{j}_{n},c^{i}_{n})\leq\epsilon_{n}$, where $\epsilon_{n}$ is the semantic radius of the category and $d$ is a distance function.
  3. 3.
    Conceptual Consistency: States within the same category share consistent characteristics such that for any $c^{1}_{n},c^{2}_{n}\in C_{n}$, $\|f(c^{1}_{n})-f(c^{2}_{n})\|\leq\delta$, where f is a feature mapping function and $\delta>0$ is a predefined threshold for similarity of features.
The Belief and CCF mechanisms work together to help an agent process environmental information and make informed decisions based on prior knowledge (Tenenbaum et al., 2011). The Belief updates beliefs through probabilistic reasoning, guiding decision-making under uncertainty. In contrast, CCF enables agents to identify and abstract conceptual features by categorizing similar experiences. CBD-RL combines these two mechanisms to make informed decisions, a unified approach (see Algorithm 1 and Theorem 4.5).
Theorem 4.5 (Cognitive Belief-Driven Reinforcement Learning).
Consider an MDP, a Conceptual Category Formation(CCF) partition $C=\{C_{n}\}_{n=1}^{N}$ of the state space $S$, and a history of experiences $H_{t}=\{(s_{i},a_{i},r_{i},s_{i+1})\}_{i=1}^{t}$. For any $s\in S$, let $\mathcal{Z}_{t}(\cdot\mid s_{t+1})$ and $\mathcal{P}_{k}(\cdot\mid s_{t})$ denote probabilities based on actions and category information, respectively, with a time-dependent adaptation parameter $\beta_{t}\in[0,1]$.
The belief-preference distribution $b_{t}:S\times A\to[0,1]$ is:
$b_{t}(\cdot\mid s_{t+1})=(1-\beta_{t})\mathcal{Z}_{t}(\cdot\mid s_{t+1})+\beta_{t}\mathcal{P}_{k}(\cdot\mid s_{t}),$ (7)
where $\beta_{t}$ is monotonic with $\lim_{t\to\infty}\beta_{t}=\beta^{*}\in[0,1]$.
4.3 CBD-RL in Discrete & Continuous Action Spaces
In the previous section, we provided a detailed explanation of the CBD-RL framework, including a brief overview, the process of Conceptual Category Formation, and the pseudocode outlining the implementation of this framework (See in Algorithm 1). However, given the diversity of experimental environments across various reinforcement learning (RL) algorithms, it is not feasible to define a single, fixed pattern for constructing our framework. Therefore, based on the nature of the action space, we have designed two distinct implementations of CBD-RL to accommodate both discrete and continuous action spaces.
Discrete action spaces consist of a finite and countable action set A, enabling explicit estimation of state-action values and precise computation of action probabilities. Upon executing an action $a_{t}\in A$, the corresponding state-action pair is assigned to its conceptual category $C_{k}$, and the action probability $\mathcal{P}_{k}(a\mid s_{t})$ is subsequently updated.
Proposition 4.6 (CBD-RL in Discrete Action Spaces).
For discrete action spaces where $A$ is finite and countable, the action selection probability for a given conceptual category $C_{k}$ is defined as:
$\mathcal{P}_{k}(a\mid s_{t})=\frac{f(a\mid s\in C_{k})}{\sum_{\tilde{a}\in A}f(\tilde{a}\mid s\in C_{k})},$ (8)
where $f(a\mid s\in C_{k})$ represents the frequency with which action $a$ has been selected in states belonging to category $C_{k}$. This probability distribution encapsulates the historical preference for actions within the category $C_{k}$.
Continuous action spaces, which optimize policy networks to maximize expected returns, present several challenges. First, decisions are based solely on the current state, limiting the use of insights from similar experiences (Haarnoja et al., 2018; Botvinick et al., 2019). Second, the value function is primarily learned via temporal-difference errors, which overlook the underlying structure of the state space. Lastly, the framework lacks mechanisms for accumulating and transferring experiential knowledge, leading to inefficiencies in long-term learning.
To overcome these limitations, we adopt probabilistic policy learning within the maximum entropy framework and extend it by incorporating category-based beliefs. Specifically, we assume the use of Gaussian distributions as the underlying probabilistic policies.
Proposition 4.7 (CBD-RL in Continuous Action Spaces).
Consider a continuous action space $A\subseteq\mathbb{R}^{n}$. For each conceptual category $C_{k}$, we maintain a Gaussian belief distribution over actions:
$\mathcal{P}_{k}(a\mid s_{t})=\mathcal{N}(\mu_{k}(s_{t}),\sigma^{2}_{k}(s_{t})),$ (9)
where $\mu_{k}(s_{t})$ and $\sigma^{2}_{k}(s_{t})$ are derived from historical transitions within category $C_{k}$ through Bayesian posterior updates. Specifically:
$\mu_{\text{posterior}}$ $=\frac{\sigma^{2}_{\text{prior}}\mu_{\text{obs}}+\sigma^{2}_{\text{obs}}\mu_{\text{prior}}}{\sigma^{2}_{\text{prior}}+\sigma^{2}_{\text{obs}}},$ (10)
$\sigma^{2}_{\text{posterior}}$ $=\frac{\sigma^{2}_{\text{prior}}\sigma^{2}_{\text{obs}}}{\sigma^{2}_{\text{prior}}+\sigma^{2}_{\text{obs}}}.$ (11)
Here, $\mu_{\text{prior}}$ and $\sigma^{2}_{\text{prior}}$ represent the mean and variance of the prior distribution, while $\mu_{\text{obs}}$ and $\sigma^{2}_{\text{obs}}$ represent the mean and variance of the observed data in the current round.

### v4

# Conceptual Belief-Informed Reinforcement Learning

@@ Abstract
Reinforcement learning (RL) has achieved significant success but is hindered by inefficiency and instability, relying on large amounts of trial-and-error data and failing to efficiently use past experiences to guide decisions. However, humans achieve remarkably efficient learning from experience, attributed to abstracting concepts and updating associated probabilistic beliefs by integrating both uncertainty and prior knowledge, as observed by cognitive science. Inspired by this, we introduce Conceptual Belief-Informed Reinforcement Learning to emulate human intelligence (HI-RL), an efficient experience utilization paradigm that can be directly integrated into existing RL frameworks. HI-RL forms concepts by extracting high-level categories of critical environmental information and then constructs adaptive concept-associated probabilistic beliefs as experience priors to guide value or policy updates. We evaluate HI-RL by integrating it into various existing value- and policy-based algorithms (DQN, PPO, SAC, and TD3) and demonstrate consistent improvements in sample efficiency and performance across both discrete and continuous control benchmarks.
@@ 1 Introduction
Reinforcement Learning (RL) has achieved remarkable success in various exciting areas, including aligning and enabling efficient inference of large language models ouyang2022training; hao2025rl, game playing (mnih2015human), robotics (singh2022reinforcement), autonomous driving (kiran2021deep), and etc. Despite these achievements, RL remains fundamentally limited by its significant sample inefficiency compared to human learning (chiu2023flexible; ye2021mastering; joshi2025shire), typically relying on vast amounts of trial-and-error interactions and often struggling to generalize to unseen or sparsely observed space (states) (mnih2015human; lake2017building). In contrast, humans can quickly learn and adapt to new spaces using only a handful of experiences, highlighting a substantial gap in data efficiency between RL and human cognition (tenenbaum2006theory; lake2015human; tenenbaum2011grow; griffiths2010probabilistic).
The gap in learning efficiency motivates the “Era of Experience” (silver2025welcome), which emphasizes leveraging past interactions to accelerate learning and foster new concepts and behaviors, rather than passively processing vast amounts of data. Cognitive science highlights two mechanisms for leveraging experience that are essential to human learning efficiency (tenenbaum2011grow): conceptual abstraction and probabilistic priors. Conceptual abstraction distills reusable structures such as prototypes, taxonomies, causal schemas — that enable compositional reasoning, generalization, and knowledge transfer (tenenbaum2011grow; lake2015human; rosch1978principles; kemp2008discovery). In parallel, behavioral studies show that humans aggregate past experiences into adaptive probabilistic priors (griffiths2005structure; peterson1967man), integrating them with future uncertainty to guide predictions and decisions (griffiths2005structure; tenenbaum2006theory).
Figure 1: Standard RL (left) replays raw transitions, while HI-RL (right) organizes them into conceptual categories with adaptive beliefs, enabling abstraction and belief-guided learning.
Various RL studies leveraged either conceptual abstraction or probabilistic priors from experience independently. However, a systematic approach to combining both experience utilization mechanisms - experience-based priors grounded in extracted conceptual formations, the very mechanism underlying humans’ efficient generalization—remains underexplored (gerstenberg2017intuitive). Specifically, abstraction in RL has focused on representation learning approaches such as contrastive learning and bisimulation metrics to compress or align observations into compact latent spaces to improve downstream task efficiency (patil2024contrastive; ferns2004metrics; castro2020scalable; peng2023conceptual). However, these methods typically do not further exploit the abstracted latent space to aggregate past experience, limiting its utility for improving RL learning efficiency. In parallel, Bayesian approaches, such as Thompson sampling, Bayesian model-based, and model-free algorithms dearden1998bayesian are widely used to address uncertainty and the exploration-exploitation tradeoff in RL (dearden1998bayesian; ghavamzadeh2015bayesian; ross2008model; thompson1933likelihood; dearden1998bayesian). However, these methods are rarely integrated with conceptual abstraction.
To efficiently leverage experience, we introduce Conceptual Belief-Informed RL, named HI-RL (Human Intelligence-RL), a framework that combines conceptual abstraction and concept-based probabilistic prior, illustrated in Figure 1. HI-RL provides an algorithm-agnostic interface that integrates seamlessly with existing RL frameworks, accelerating learning by leveraging experience efficiently. It extracts concepts from large state spaces and reformulates experience into priors grounded in these abstractions, mimicking human-like conceptualization for learning efficiency. Our main contributions are summarized as below:
  • We present HI-RL , an experience-utilization framework that efficiently leverages past experiences to emulate human-like learning efficiency. HI-RL reformulates the set of past experiences into probabilistic belief priors grounded in conceptual abstractions. These concept-based priors are adaptively updated over time and incorporated as auxiliary knowledge into RL value or policy updates.
  • HI-RL is algorithm-agnostic and functions as a flexible module that can be seamlessly integrated into existing RL frameworks. To demonstrate its versatility, we integrate HI-RL into several popular RL algorithms (Q-learning, PPO, SAC, and TD3) and evaluate performance across both discrete and continuous tasks, achieving consistent improvements in learning efficiency and overall performance.
@@ 2 Related Works
2.1 Cognitive Science for Conceptual Learning
@@
Humans achieve remarkable learning efficiency by generalizing from limited experience through Bayesian inference, integrating prior knowledge with new evidence under uncertainty (tenenbaum2001generalization; griffiths2005structure; tenenbaum2006theory). This supports conceptual abstraction—extracting high-level structure from sparse data—and enables causal reasoning and cross-domain transfer (tenenbaum2011grow; kemp2008discovery). Recent work formalizes how learners reorganize internal knowledge via probabilistic reasoning (lake2015human; lake2017building), motivating the integration of such principles into machine learning for scalability, adaptability, and sample efficiency (ma2022learning). Studies further show that uncovering latent causal structures enhances interpretability and abstraction, even in complex domains such as joint behavior analysis (gu2025causkelnet; gu2024advancing). Yet, despite these advances, reinforcement learning remains dominated by replay, metric-based similarity, or policy integration, with little use of structured conceptual abstraction from cognitive science.
2.2 Experience-Informed Reinforcement Learning
Experience has long been exploited to improve efficiency in RL. Habit-based RL models long-term regularities as habitual priors that accelerate action selection but lack flexibility for abstraction and transfer (daw2005uncertainty; collinsbeyond; keramati2011speed). Replay-based techniques such as PER (schaul2015prioritized), HER and its prioritized variants (andrychowicz2017hindsight; sun2025hierarchical; kim2025extended) enhance sample efficiency by weighting or relabeling transitions, while refinements like FoDA (chen2024foresight) and EDER (zhao2024efficient) adapt distributions or promote diversity to improve generalization. Beyond replay, episodic memory models (NEC) (pritzel2017neural) enable rapid value retrieval, and hybrid gradients (Q-Prop, IPG) (gu2016q; gu2017interpolated) fuse on- and off-policy signals for variance reduction. Collectively, these methods leverage past interactions via sampling or memory mechanisms, yet remain confined to buffer-level operations and lack pathways for higher-order conceptual abstraction and belief-structured generalization.
2.3 Abstraction in Reinforcement Learning
State abstraction has long been studied as a means to compress state spaces and enable generalization in RL (bertsekas1988adaptive; givan2003equivalence; ravindran2004algebraic; ravindran2003smdp; li2006towards; kulkarni2016hierarchical). Classical bisimulation and Kantorovich metrics provide strong theoretical guarantees but are computationally expensive and highly sensitive to perturbations (ferns2004metrics; ferns2011bisimulation). Task-specific metrics improve offline evaluation (pavse2023state) but lack adaptability, while scalable relaxations (castro2020scalable) trade rigor for tractability. Trajectory-chain and pseudometric methods (girgin2007state; dadashi2021offline) offer finer granularity but incur high storage or auxiliary costs. More recent work, such as (patil2024contrastive), leverages contrastive objectives and modern Hopfield networks to compress large state spaces into abstract nodes, thereby facilitating downstream RL. These approaches primarily focus on constructing a new, compressed state space or representation for downstream algorithms. In contrast, our framework preserves the original state and exploration space while introducing an abstraction-based belief layer on top. We focus on utilizing conceptual abstraction as a basis to update its probabilistic priors, efficiently aggregating and using past experience to improve toward human-like efficient learning.
@@ 3 Problem Formulation
Markov Decision Process (MDP) Considering reinforcement learing problems formalized as MDP (bellman1957markovian; sutton2018reinforcement) $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{T},r,\mu_{0},\gamma,T)$. Here $T$ is the horizon length. $\mathcal{S}$ denotes states space ($s\in S$) and $\mathcal{A}$ denotes the action spaces ($a\in A$). $\mathcal{T}(s_{t+1}\mid s_{t},a_{t})$ represents the transition dynamics, specifying the probability distribution over the next state $s_{t+1}$ conditioned on the current state $s_{t}\in\mathcal{S}$ and action $a_{t}\in\mathcal{A}$ at $t_{th}$ time step $(1\leq t\leq T)$. $r(s,a)$ represents the reward function given the state s and action a. The initial state follows $\mu_{0}$, $\gamma\in(0,1)$ is the discount factor. The goal of this MDP problem is to identify an optimal policy $\pi$ that achieves the maximum expected discounted return:
$\max_{\pi}\;\mathbb{E}_{\pi,\mathcal{T},\mu_{0}}\Big[\sum_{t=0}^{T}\gamma^{t}r(s_{t},a_{t})\Big].$ (1)
Formally, given the horizon length $T$ and transition dynamics $\mathcal{T}$, the long-term return from time step $t=0$ to $t=T$ associated with the optimal policy $\pi$ is quantified through the Q-function and the Value-function (watkins1992q) by expected cumulative rewards from initial state $\mu_{0}$, defined as:
$Q^{\pi}(s,a)=\mathbb{E}_{\pi,\mathcal{T},\mu_{0}}\Big[\sum_{t=0}^{T}\gamma^{t}r(s_{t},a_{t})\;\big|\;s_{0}=s,a_{0}=a\Big],V^{\pi}(s)=\mathbb{E}_{\pi,\mathcal{T},\mu_{0}}\Big[\sum_{t=0}^{T}\gamma^{t}r(s_{t},a_{t})\;\big|\;s_{0}=s\Big]$ (2)
4 Conceptual Belief-Informed Reinforcement Learning
In this section, we present Conceptual Belief-Informed Reinforcement Learning (HI-RL ), enhancing experience to emulate human intelligence learning efficiency. HI-RL consists of two core modules: (i) Concept Formation, which clusters state–based experiences into semantically coherent categories, and (ii) a belief representing probabilistic prior grounded on different concepts, defining probabilistic action experience prior over these categories. By coupling conceptual abstraction with belief-guided reasoning, HI-RL provides a structured and uncertainty-aware foundation for policy learning, supporting stable updates, efficient generalization, and reuse of past experiences.
4.1 Concept Formation
The foundation for abstracting concepts can vary, as long as it represents the current situation and critical information about the environment and the agent. In this work, HI-RL focuses on state spaces, as states encapsulate essential information for decision-making and directly influence the agent’s behavior and learning process, enabling pattern recognition and generalization. Specifically, we partition the states into disjoint subsets, with each subset representing a distinct concept formed by grouping states with shared characteristics and properties, thereby facilitating effective knowledge transfer within each concept. In the following, we mathematically define a conceptual abstraction as:
Definition 4.1 (Concept Formation in State Space).
A concept formation in the state space is defined as a collection of subsets $C_{K}=\{C_{1},\dots,C_{K}\}$ that satisfy $\mathcal{S}=\bigcup_{k=1}^{K}C_{k}$, meaning the subsets are disjoint and collectively cover the entire state space. Here, $K$ denotes a finite, prescribed number of concept categories.
In practice, conceptual abstractions can be obtained with various clustering methods. In this work, we adopt K-means (lloyd1982least) for its simplicity and scalability, though alternatives (e.g., spectral or hierarchical clustering) are equally applicable.
4.2 Conceptual Adaptive Belief for RL
With abstract concepts in mind, where each concept groups states that share similar features and actions, we aggregate observed information within a concept into a unified container, a concept-based belief. Philosophically, a belief is an internal representation of how an agent interprets and anticipates the world, serving as a guide for inference and decision-making under uncertainty rather than as absolute truth (dennett1988precis). In this work, each concept $C_{k}$ is paired with a time-adaptive belief $b_{t}(\cdot\mid k)\in\Delta(\mathcal{A})$, derived from the accumulation of past decisions and outcomes within that concept. Formally, for a conceptual abstraction $C_{K}=\{C_{1},\dots,C_{K}\}$, we define the mapping $b_{t}:[K]\to\Delta(\mathcal{A})$, where $b_{t}(\cdot\mid k)$ encodes the integrated action preferences of all states belonging to $C_{k}$.
We leverage the aggregated experience within each concept to accelerate learning by using concept-based beliefs as priors in RL updates. These beliefs can be seamlessly integrated into any existing RL algorithm. At each timestep $t$, we combine two signals: (i) instant feedback $\mathcal{Z}_{t}:\mathcal{S}\to\mathcal{A}$, defined by the base algorithm (e.g., Q-values in DQN, Gaussian policy in SAC, clipped surrogate in PPO, or deterministic actor in TD3), and (ii) the prior $b_{t}$, aggregated from past experience within the corresponding concept. For a given state $s\in\mathcal{S}$, we first identify its concept index $c(s)$ such that $s\in C_{c(s)}$, and then fuse the signals as:
$B_{t}(\cdot\mid s)=(1-\beta_{t})\mathcal{Z}_{t}(\cdot\mid s)+\beta_{t}b_{t}(\cdot\,|\,c(s)),$ (3)
where $\beta_{t}\in[0,1]$ is an adaptive parameter monotonic in $t$, satisfying $\lim_{t\to\infty}\beta_{t}=\beta^{*}$ with $\beta^{*}\in[0,1]$ a constant denoting the limiting weight on conceptual priors. In this formulation, $b_{t}(\cdot\mid c(s))$ is the empirical concept-based prior aggregated from experience, while $B_{t}(\cdot\mid s)$ is the fused distribution actually used for decision-making by combining $b_{t}$ with the instant feedback $\mathcal{Z}_{t}$.
This formulation ensures that the decision-making solutions for every state $s_{t}$ are influenced by both immediate feedback from the environment and the prior experience derived from the conceptual abstraction $C_{c(s)}$ to which it belongs.
Algorithm 1 Conceptual Belief-Informed RL (HI-RL)
1:Initialize concept priors $b(\cdot\mid c(s))$
2:for $t=1,2,\dots$ do
3:  Observe $s_{t}$; form $\mathcal{Z}_{t}(\cdot\mid s_{t})$; fuse $B_{t}(\cdot\mid s_{t})=(1-\beta_{t})\mathcal{Z}_{t}(\cdot\mid s_{t})+\beta_{t}b_{t}(\cdot\mid c(s_{t}))$
4:  Sample $a_{t}\!\sim\!B_{t}$; step env $(r_{t},s_{t+1})$
5:  Update policy with $B_{t}$ and prior experience $b_{t}$
6:end for